Resistors for electronic circuits are manufactured on a high-speed automated machine. The machine is set up to produce a large run of resistors of 1,000 ohms each. Use Exhibit 10.13.

     To set up the machine and to create a control chart to be used throughout the run, 15 samples were taken with four resistors in each sample. The complete list of samples and their measured values are as follows: Use three-sigma control limits.

SAMPLE NUMBER READINGS (IN OHMS)
1 1000 992 1014 1026
2 1006 977 1027 1016
3 985 1029 1018 992
4 1001 1006 971 1000
5 1028 970 979 1025
6 997 1026 1007 1003
7 1016 1003 987 1006
8 995 1029 978 986
9 1027 995 1003 1008
10 991 987 981 996
11 994 1000 978 971
12 999 983 973 1005
13 1003 993 974 1002
14 1005 972 994 1000
15 1012 1018 982 1014

a. Calculate the mean and range for the above samples. (Round "Mean" to 2 decimal places and "Range" to the nearest whole number.)

Sample Number Mean Range
1
2
3
4
5
6
7
8
9
10
11
12
13
14
15

b. Determine X=X= and R−R−. (Round your answers to 3 decimal places.)

X=X=
R−R−

c. Determine the UCL and LCL for a X−X−chart. (Round your answers to 3 decimal places.)

UCL
LCL

d. Determine the UCL and LCL for R-chart. (Leave no cells blank - be certain to enter "0" wherever required. Round your answers to 3 decimal places.)

UCL
LCL

e. What comments can you make about the process?

The process is in statistical control.
The process is out of statistical control.
Answer #1

a.

Sample Mean and Sample Range for all the samples are calculated as per the below table:

Sample Readings
Number 1 2 3 4 Sample Mean Sample Range
1 1000 992 1014 1026 1008 34
2 1006 977 1027 1016 1006.5 50
3 985 1029 1018 992 1006 44
4 1001 1006 971 1000 994.5 35
5 1028 970 979 1025 1000.5 58
6 997 1026 1007 1003 1008.25 29
7 1016 1003 987 1006 1003 29
8 995 1029 978 986 997 51
9 1027 995 1003 1008 1008.25 32
10 991 987 981 996 988.75 15
11 994 1000 978 971 985.75 29
12 999 983 973 1005 990 32
13 1003 993 974 1002 993 29
14 1005 972 994 1000 992.75 33
15 1012 1018 982 1014 1006.5 36

b.

X double bar = Mean of Sample Ranges = ( 1008 + 1006.5 + 1006 + 994.5 + 1000.5 + 1008.25 + 1003 + 997 + 1008.25 + 988.75 + 985.75 + 990 + 993 + 992.75 + 1006.5) / 15 = 14988.75 / 15 = 999.25

R bar = Mean of Sample Rnages = ( 34 + 50 + 44 + 35 + 58 + 29 + 29 + 51 + 32 + 15 + 29 + 32 + 29 + 33 + 36) / 15 = 536 / 15 = 35.733

c.

X bar chart:

UCL = X double bar + ( R bar x A2)

LCL = X double bar - ( R bar x A2)

For a sample size of 4, A2 value is 0.729 ( From Control chart constant table)

So,

UCL = 999.25 + ( 35.733 x 0.729) = 999.25 + 26.049 = 1025.299

LCL = 999.25 - ( 35.733 x 0.729) = 999.25 - 26.049 = 973.201

d.

R bar chart:

UCL = R bar x D4

LCL = R bar x D3

For a sample size of 4, D4 value is 2.282 and D3 value is 0 ( From Control chart constant table)

So,

UCL = 35.733 x 2.282 = 81.543

LCL = 35.733 x 0 = 0

e.

All the Sample mean values are within the UCL and LCL values of X bar chart.

All the Sample range values are within the UCL and LCL values of R bar chart.

Therefore,

The process is in statistical control.